# Hierarchical Large Scale Multirobot Path (Re)Planning

### v1

@@ VIII-B Effectiveness of High-level Planning
To demonstrate the effectiveness of our high-level planner, we compare the cell congestion of the proposed MCF-based algorithm to an egocentric greedy approach. The greedy planner outputs a single-robot-based shortest inter-cell routing without considering other robots’ routing, and may lead to congestion in certain partitions. Additionally, we compare the MCF-based approach with the baseline by imposing the partition into the workspace.
In Table. I, we report the average high-level computation time $\bar{t}_{\mathrm{high}}$ and the averaged maximum number of robots in a cell throughout the whole execution $\bar{N}_{\mathrm{max}}$. The complexity of a MAPF instance scales poorly with the number of robots. Thus, $\bar{N}_{\mathrm{max}}$ is a good indicator of the computational hardness of a MAPF instance. The computation time for greedy planning is instant while the MCF-based methods have additional overhead that increases with the number of cells, since they are centralized. For all instances in this paper, MCF-based methods provide real-time solutions.
The MCF-based methods reduce the congestion compared to greedy and baseline methods, according to $N_{\mathrm{max}}$. Although the advantage to the baseline is minor in the current experiment setup, the MCF-based approach can further reduce the congestion by relaxing bounded-suboptimality constraints, i.e., increase $w_{mcf}$. For a fair comparison of low-level planning time, both greedy and MCF-based utilize the same number of threads in computation. Compared to the greedy approach, the proposed MCF-based high-level planning yields more efficient low-level planning, which brings computation time to real time while maintaining the solution quality, i.e., makespan.
@@ VIII-C Scalability of the Proposed Algorithm
To investigate the scalability of the proposed algorithm, we run simulations on increasing numbers of robots (c.f., Table. I). Note that the proposed algorithm achieves real-time performance in all instances as we scale up the number of robots, by looking at the averaged maximum high-level and low-level replanning time $\bar{t}^{\mathrm{max}}_{\mathrm{high}}$ and $\bar{t}^{\mathrm{max}}_{\mathrm{low}}$. Due to the hierarchical approach, the proposed algorithm can further scale to larger teams with more cells and CPU threads. The computation bottleneck is the centralized high-level planning in larger scale problems, which we aim to address in future work.
@@ VIII-D Physical Robots
Fig. 1 shows a representative experiment with $32$ physical Crazyflies in a cluttered environment (video link: https://youtu.be/ftdWVpLkErs). We use a Vicon motion capture system to localize and CrazySwarm [23] to control the robots. In the experiment, initial positions are uniformly spaced on an ellipse within the workspace, and the goal positions are the antipodal points on the ellipse. Three column obstacles are placed within the ellipse. The experiment demonstrates that the proposed algorithm distributes robots effectively through the workspace and achieves real-time replanning.
@@ IX Conclusion and Future Work
We have introduced a hierarchical path planning algorithm for large-scale coordination tasks. Our algorithm significantly reduces the computation time and suits on-demand applications such as drone delivery despite yielding a suboptimal solution. The framework achieves real-time operation by dividing the workspace into disjoint convex cells; within each, an anytime MAPF planner computes collision-free paths in parallel. Our high-level planner regulates the congestion of regions while guaranteeing the routing quality. Additionally, our algorithm considers the robot’s geometric shape constraints in continuous space, and we run experiments with the collision model of a quadrotor with downwash. We also devise a cell-crossing protocol, which guarantees the robot always has a plan, even when transiting between cells, and allows replanning in a continuous time domain.
The proposed algorithm is designed for lifelong replanning. In our experiments, goals are chosen from a pre-determined set, however, it can be extended to a general lifelong replanning as we add and delete goals online. When new goals are requested, the algorithm can identify the new goal’s cell, and run conflict annotation steps in parallel for each new goal.
While the MCF-based high-level planner operates in real-time in our experiments with up to $142$ robots, the limits of its real-time operation are not well defined and depend on the number and density of robots and cells in the space. Future work will explore distributed MCF [24, 25] to increase scalability of the system with real-time, distributed operation regardless of the number of robots and cells. Furthermore, we aim to solve real time large scale motion planning, which respects the robot’s kinodynamic constraints and plans in continuous space and time.
X Acknowledgements
The authors would like to thank Baskın Şenbaşlar, Eric Ewing, Yuan Yuan, Calvin Luo, Yutong Wang, Arjun Prakash for their help in revising the manuscript.

### v2

@@ VIII-B Effectiveness of High-level Planning
To demonstrate the effectiveness of our MCF-based high-level planner, we compare its cell congestion to an egocentric greedy approach. The greedy planner outputs a single-robot-based shortest inter-cell routing without considering other robots’ routing, and may lead to congestion in certain cells. Additionally, we compare the MCF-based approach with the baseline by imposing the partition onto the workspace.
In Table I, we report the average high-level computation time $\bar{t}_{\mathrm{high}}$ and the averaged maximum number of robots in a cell throughout the whole execution $\bar{N}_{\mathrm{max}}$. The complexity of a MAPF instance scales poorly with the number of robots. Thus, $\bar{N}_{\mathrm{max}}$ is an insightful indicator of the computational hardness of a MAPF instance. The computation time for greedy high-level planning is instant, while the MCF-based methods have additional overhead that increases with the number of cells, since they are centralized. For all instances in this paper, MCF-based methods provide real-time solutions.
The MCF-based methods reduce the congestion compared to greedy and baseline methods, according to $N_{\mathrm{max}}$. Although the advantage to the baseline is minor in the current experiment setup, the MCF-based approach can further reduce the congestion by relaxing the suboptimality bound, i.e., increase $w_{mcf}$. For a fair comparison of low-level planning time, both greedy and MCF-based utilize the same number of threads in computation. Compared to the greedy approach, the proposed MCF-based inter-cell routing brings cell planning computation time to real time while maintaining its solution quality, i.e., makespan.
@@ VIII-C Scalability of the Proposed Algorithm
To validate the scalability of our algorithm, we run simulations on increasing numbers of robots (c.f., Table. I). Note that our algorithm achieves real-time performance in all instances as we scale up, according to the averaged maximum high-level and low-level replanning time $\bar{t}^{\mathrm{max}}_{\mathrm{high}}$ and $\bar{t}^{\mathrm{max}}_{\mathrm{low}}$. Due to our hierarchical approach, it can further scale to larger teams with more cells and CPU threads. The computation bottleneck is the centralized high-level planning in larger scale problems, which we aim to address in future work.
@@ VIII-D Physical Robots
Fig. 1 shows a representative experiment with 32 physical Crazyflies in a cluttered environment (video link: https://youtu.be/ftdWVpLkErs). We use a Vicon motion capture system to localize and CrazySwarm [25] to control the robots. In the experiment, robots are uniformly located on an ellipse, with their antipodal goals. Three column obstacles are placed within the ellipse. The experiment demonstrates that the proposed algorithm distributes robots effectively through the workspace and achieves real-time replanning.
@@ IX Conclusion and Future Work
We have introduced a hierarchical path planning algorithm for large-scale coordination tasks. Despite yielding a suboptimal solution, our algorithm significantly reduces the computation time and suits on-demand applications such as drone delivery. The framework achieves real-time operation by dividing the workspace into disjoint cells; within each, an anytime MAPF planner computes collision-free paths in parallel. Our high-level planner regulates congestion while guaranteeing routing quality. Additionally, our algorithm considers the robot embodiment, and we run experiments with the downwash-aware collision model of a quadrotor. We also devise a cell-crossing protocol, which guarantees non-stop execution even when transiting between cells.
The proposed algorithm is designed for lifelong replanning. In our experiments, goals are chosen from a pre-determined set, however, it can be extended to a lifelong replanning as we request new goals online, and the algorithm runs conflict annotation to update the conflicts.
While the MCF-based high-level planner operates in real-time in our experiments with up to 142 robots in a 25-cell partition, the limits of its real-time operation are not well defined and depend on the number and density of robots and cells in the space. Future work will explore distributed MCF [26, 27] to increase scalability of the system with real-time, distributed operation regardless of the number of robots and cells. Furthermore, we aim to solve real time large scale motion planning, which respects the robot’s kinodynamic constraints and plans in continuous space and time.